Home‎ > ‎Crew 143‎ > ‎

1124-CommandersReport

Commander Report
11/24/2014
Paul Knightly

The crew has been wrapping up their scientific work to perform on EVA the past couple of days as the science portion of the mission begins to wind down. Four of us went EVA today to explore additional targets to the east and fresh targets to the west of the Hab. Paul Sokoloff finished collecting plant and other biological samples for his study and the group, including Claude-Michel, Anastasiya and myself conducted another soil sampling test for Alexandre’s SSLP experiment.
We finally made our way to the top of the plateau west of the Hab and the view was absolutely breathtaking. We spent the remainder of the EVA collected rock and plant samples and Anastasiya continued her search for meteorites but to no avail. All good things must come to an end and the EVA concluded just before noon.
The crew spent the remainder of the afternoon post-processing samples and pictures from the EVA. Paul S. provided another first aid briefing discussing proper procedures for handling broken bones.
Today is the start of our second shower cycle at the Hab and some time was spent getting the water heater to start up – it seems to be functioning normally now. We also trapped a stowaway mouse in the galley overnight, bringing our mission tally to two. Fortunately, we have been regularly sterilizing all food preparation and eating surfaces and do not anticipate any ill-affects from our rodent friends.
We are gearing up for construction activities on the tunnel to the observatory and are tentatively planning a short EVA tomorrow to test out construction techniques before diving into a full construction EVA. Otherwise, tomorrow will be a lab day for the crew to organize and catch up on work as our time on Mars draws to a close.